L^{A}TEX 2ε for authors

\begin{matrix}c \\ ◯ \end{matrix} Copyright 1994, L^{A}TEX3 Project Team. All rights reserved.

14 December 1994

Contents

1 Introduction 2 1.1 L^{A}TEX 2ε—The new L^{A}TEX release . . . . . . . . . . . . . . . . . . 2 1.2 L^{A}TEX3—The long-term future of L^{A}TEX . . . . . . . . . . . . . . 3 1.3 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.4 Further information . . . . . . . . . . . . . . . . . . . . . . . . . 4

2 Classes and packages 4 2.1 What are classes and packages? . . . . . . . . . . . . . . . . . . . 4 2.2 Class and package options . . . . . . . . . . . . . . . . . . . . . . 6 2.3 Standard classes . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.4 Standard packages . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.5 Related software . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

3 Commands 10 3.1 Initial commands . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.2 Preamble commands . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.3 Document structure . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.4 Definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.5 Boxes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.6 Measuring things . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.7 Line endings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.8 Controlling page breaks . . . . . . . . . . . . . . . . . . . . . . . 17 3.9 Floats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.10 Font changing: text . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.11 Font changing: math . . . . . . . . . . . . . . . . . . . . . . . . . 19 3.12 Ensuring math mode . . . . . . . . . . . . . . . . . . . . . . . . . 19 3.13 Text commands: all encodings . . . . . . . . . . . . . . . . . . . 20 3.14 Text commands: the T1 encoding . . . . . . . . . . . . . . . . . . 22 3.15 Logos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 3.16 Picture commands . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.17 Old commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 4 L^{A}TEX 2.09 documents 24

4.1 Font selection problems . . . . . . . . . . . . . . . . . . . . . . . 24 4.2 Native mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

5 Problems 25 5.1 New error messages . . . . . . . . . . . . . . . . . . . . . . . . . . 26 5.2 Old internal commands . . . . . . . . . . . . . . . . . . . . . . . 28 5.3 Old files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 5.4 Where to go for more help . . . . . . . . . . . . . . . . . . . . . . 29

6 Enjoy! 29

1 Introduction

Welcome to L^{A}TEX 2ε, the new standard version of the L^{A}TEX Document Preparation System.

This document describes how to take advantage of the new features of L^{A}TEX, and how to process your old L^{A}TEX documents with L^{A}TEX 2ε. However, this document is only a brief introduction to the new facilities and is intended for authors who are already familiar with the old version of L^{A}TEX. It is not a reference manual for L^{A}TEX 2ε.

1.1 L^{A}TEX 2ε—The new L^{A}TEX release

The previous version of L^{A}TEX was known as L^{A}TEX 2.09. Over the years many extensions have been developed for L^{A}TEX. This is, of course, a sure sign of its continuing popularity but it has had one unfortunate result: incompatible L^{A}TEX formats came into use at different sites. This included ‘standard L^{A}TEX 2.09’, L^{A}TEX built with the New Font Selection Scheme (NFSS), SliTEX, AMS-L^{A}TEX, and so on. Thus, to process documents from various places, a site maintainer was forced to keep multiple versions of the L^{A}TEX program. In addition, when looking at a source file it was not always clear for which format the document was written.

To put an end to this unsatisfactory situation, L^{A}TEX 2ε has been produced; it brings all such extensions back under a single format and thus prevents the proliferation of mutually incompatible dialects of L^{A}TEX 2.09. With L^{A}TEX 2ε the ‘new font selection scheme’ is standard and, for example, amstex (formerly the AMS-L^{A}TEX format) or slides (formerly the SliTEX format) are simply extension packages, all working with the same base format.

The introduction of a new release also made it possible to add a small number of often-requested features and to make the task of writing packages and classes simpler.

1.2 L^{A}TEX3—The long-term future of L^{A}TEX

L^{A}TEX 2ε is the consolidation step in a comprehensive reimplementation of the L^{A}TEX system. The next major release of L^{A}TEX will be L^{A}TEX3, which will include a radical overhaul of the document designers’ and package writers’ interface to L^{A}TEX.

L^{A}TEX3 is a long-term research project but, until it is completed, the project team are committed to the active maintenance of L^{A}TEX 2ε. Thus the experience gained from the production and maintenance of L^{A}TEX 2ε will be a major influence on the design of L^{A}TEX3.

If you would like to support the project then you are welcome to send donations to the L^{A}TEX3 Project Fund; this has been set up to help the research team by financing various expenses associated with this voluntary work of maintaining the current L^{A}TEX and developing L^{A}TEX3.

The fund is administered by The TEX Users Group and by various local user groups. Information about making donations and joining these groups is available from:

TEX Users Group

P. O. Box 869, Santa Barbara, CA 93102-0869, USA

Fax: +1 805 963 8358

     E-Mail: tug@tug.org

Articles covering aspects of the L^{A}TEX3 project are available by anonymous ftp in the Comprehensive TEX Archive, in the directory: tex-archive/info/ltx3pub

The file ltx3pub.bib in that directory contains an abstract of each of the files.

1.3 Overview

This document contains an overview of the new structure and features of L^{A}TEX. It is not a self-contained document, as it contains only the features of L^{A}TEX which have changed since version 2.09. You should read this document in conjunction with an introduction to L^{A}TEX.

Section 2 contains an overview of the new structure of L^{A}TEX documents. It describes how classes and packages work and how class and package options can be used. It lists the standard packages and classes which come with L^{A}TEX.

Section 3 describes the new commands available to authors in L^{A}TEX 2ε.

Section 4 shows how to process old L^{A}TEX documents with L^{A}TEX 2ε.

Section 5 contains advice on dealing with problems you may encounter in running L^{A}TEX 2ε. It lists some error messages which are new in L^{A}TEX 2ε and it describes some of the more common problems and how to cure them, or where to find further information.

1.4 Further information

For a general introduction to L^{A}TEX, including the new features of L^{A}TEX 2ε, you should read L^{A}TEX: A Document Preparation System by Leslie Lamport [3].

A more detailed description of the new features of L^{A}TEX, including an overview of more than 150 packages, is to be found in The L^{A}TEX Companion by Michel Goossens, Frank Mittelbach and Alexander Samarin [1].

For more information about the many new L^{A}TEX packages you should read the package documentation, which should be available from the same source as your copy of L^{A}TEX.

There are a number of documentation files which accompany every copy of L^{A}TEX. A copy of L^{A}TEX News will come out with each six-monthly release of L^{A}TEX; it will be found in the files ltnews*.tex. The class- and package-writer’s guide L^{A}TEX 2ε for Class and Package Writers describes the new L^{A}TEX features for writers of document classes and packages; it is in clsguide.tex. The guide L^{A}TEX 2ε Font Selection describes the L^{A}TEX font selection scheme for class- and package-writers; it is in fntguide.tex.

We are gradually turning the source code for L^{A}TEX into a L^{A}TEX document L^{A}TEX: the program. This document includes an index of L^{A}TEX commands and can be typeset from source2e.tex.

For more information about TEX and L^{A}TEX, please contact your local TEX Users Group, or the international TEX Users Group, P. O. Box 869, Santa Barbara, CA 93102-0869, USA, Fax: +1 805 963 8358, E-mail: tug@tug.org.

2 Classes and packages

This section describes the new structure of L^{A}TEX documents and the new types of file: classes and packages.

2.1 What are classes and packages?

The main difference between L^{A}TEX 2.09 and L^{A}TEX 2ε is in the commands before

\begin{document}.

In L^{A}TEX 2.09, documents had styles, such as article or book, and options, such as twoside or epsfig. These were indicated by the \documentstyle command:

     \documentstyle[⟨options⟩]{⟨style⟩}

For example, to specify a two-sided article with encapsulated PostScript figures, you said:

   \documentstyle[twoside,epsfig]{article}

However, there were two different types of document style option: built-in options such as twoside; and packages such as epsfig.sty. These were very different, since any L^{A}TEX document style could use the epsfig package but only document styles which declared the twoside option could use it.

To avoid this confusion, L^{A}TEX 2ε differentiates between built-in options and packages. These are given by the new \documentclass and \usepackage commands:

     \documentclass[⟨options⟩]{⟨class⟩}

\usepackage[⟨options⟩]{⟨packages⟩}

For example, to specify a two-sided article with encapsulated PostScript figures, you now write:

   \documentclass[twoside]{article}
   \usepackage{epsfig}

You can load more than one package with a single \usepackage command; for example, rather than writing:

   \usepackage{epsfig}
   \usepackage{multicol}

you can specify:

   \usepackage{epsfig,multicol}

Note that L^{A}TEX 2ε still understands the L^{A}TEX 2.09 \documentstyle command. This command causes L^{A}TEX 2ε to enter L^{A}TEX 2.09 compatibility mode, which is described in Section 4.

You should not, however, use the \documentstyle command for new documents because this compatibility mode is very slow and the new features of L^{A}TEX 2ε are not available in this mode.

To help differentiate between classes and packages, document classes now end with .cls rather than .sty. Packages still end with .sty, since most L^{A}TEX 2.09 packages work well with L^{A}TEX 2ε.

2.2 Class and package options

In L^{A}TEX 2.09, only document styles could have options such as twoside or draft. In L^{A}TEX 2ε, both classes and packages are allowed to have options. For example, to specify a two-sided article with graphics using the dvips driver, you write:

   \documentclass[twoside]{article}
   \usepackage[dvips]{graphics}

It is possible for packages to share common options. For example, you could, in addition, load the color package by specifying:

   \documentclass[twoside]{article}
   \usepackage[dvips]{graphics}
   \usepackage[dvips]{color}

But because \usepackage allows more than one package to be listed, this can be shortened to:

   \documentclass[twoside]{article}
   \usepackage[dvips]{graphics,color}

In addition, packages will also use each option given to the \documentclass command (if they know what to do with it), so you could also write:

   \documentclass[twoside,dvips]{article}
   \usepackage{graphics,color}

Class and package options are covered in more detail in The L^{A}TEX Companion and in L^{A}TEX 2ε for Class and Package Writers.

2.3 Standard classes

The following classes are distributed with L^{A}TEX:

article The article class described in L^{A}TEX: A Document Preparation System.

book The book class described in L^{A}TEX: A Document Preparation System.

report The report class described in L^{A}TEX: A Document Preparation System.

letter The letter class described in L^{A}TEX: A Document Preparation System. slides The slides class described in L^{A}TEX: A Document Preparation System,

formerly SliTEX.

proc A document class for proceedings, based on article. Formerly the proc package.

ltxdoc The document class for documenting the L^{A}TEX program, based on

article.

ltxguide The document class for L^{A}TEX 2ε for Authors and L^{A}TEX 2ε for Class and Package Writers, based on article. The document you are reading now uses the ltxguide class. The layout for this class is likely to change in future releases of L^{A}TEX.

ltnews The document class for the L^{A}TEX News information sheet, based on article. The layout for this class is likely to change in future releases of L^{A}TEX.

2.4 Standard packages

The following packages are distributed with L^{A}TEX:

alltt This package provides the alltt environment, which is like the verbatim

New feature 1994/12/01

environment except that \, {, and } have their usual meanings. It is described in alltt.dtx and L^{A}TEX: A Document Preparation System.

doc This is the basic package for typesetting the documentation of L^{A}TEX programs. It is described in doc.dtx and in The L^{A}TEX Companion.

exscale This provides scaled versions of the math extension font. It is described

in exscale.dtx.

fontenc This is used to specify which font encoding L^{A}TEX should use. It is

described in ltoutenc.dtx.

graphpap This package defines the \graphpaper command; this can be used

New feature 1994/12/01

in a picture environment.

ifthen Provides commands of the form ‘if… then do… otherwise do… ’. It is described in ifthen.dtx and The L^{A}TEX Companion.

inputenc This is used to specify which input encoding L^{A}TEX should use. It is

New feature 1994/12/01

described in inputenc.dtx.

latexsym L^{A}TEX 2ε no longer loads the L^{A}TEX symbol font by default. To access it, you should use the latexsym package. It is described in latexsym.dtx and in The L^{A}TEX Companion; see also Section 5. makeidx This provides commands for producing indexes. It is described in

L^{A}TEX: A Document Preparation System and in The L^{A}TEX Companion.

newlfont This is used to emulate the font commands of L^{A}TEX 2.09 with the New Font Selection Scheme. It is described in The L^{A}TEX Companion.

oldlfont This is used to emulate the font commands of L^{A}TEX 2.09. It is described in The L^{A}TEX Companion.

showidx This causes the argument of each \index command to be printed on the page where it occurs. It is described in L^{A}TEX: A Document Preparation System.

syntonly This is used to process a document without typesetting it. It is described in syntonly.dtx and in The L^{A}TEX Companion.

tracefnt This allows you to control how much information about L^{A}TEX’s font loading is displayed. It is described in The L^{A}TEX Companion.

2.5 Related software

The following software should be available from the same distributor as your copy of L^{A}TEX 2ε. You should obtain at least the graphics and tools collections in order to have all the files described in L^{A}TEX: A Document Preparation System. The amstex package (part of amslatex) and babel are also mentioned in the list of ‘standard packages’ in section C.5.2 of that book.

amslatex Advanced mathematical typesetting from the American Mathematical Society. This is includes the amstex package; it provides many commands for typesetting mathematical formulas of higher complexity. It is produced and supported by the American Mathematical Society and it is described in The L^{A}TEX Companion.

babel This package and related files support typesetting in many languages. It is described in The L^{A}TEX Companion.

graphics This includes the graphics package which provides support for the inclusion and transformation of graphics, including files produced by other software. Also included, is the color package which provides support for typesetting in colour. Both these packages are described in L^{A}TEX: A Document Preparation System.

mfnfss Everything you need (except the fonts themselves) for typesetting with a large range of bit-map (Metafont) fonts.

psnfss Everything you need (except the fonts themselves) for typesetting with a large range of Type 1 (PostScript) fonts. tools Miscellaneous packages written by the L^{A}TEX3 project team. These packages come with documentation and each of them is also described in at least one of the books The L^{A}TEX Companion and L^{A}TEX: A Document Preparation System.

2.5.1 Tools This collection of packages includes, at least, the following (some files may have slightly different names on certain systems):

array Extended versions of the environments array, tabular and tabular*,

with many extra features. dcolumn Alignment on ‘decimal points’ in tabular entries. Requires the array

package. delarray Adds ‘large delimiters’ around arrays. Requires array. hhline Finer control over horizontal rules in tables. Requires array. longtable Multi-page tables. (Does not require array, but it uses the extended

features if both are loaded.) tabularx Defines a tabularx environment that is similar to tabular* but it

modifies the column widths, rather than the inter-column space, to achieve the desired table width.

afterpage Place text after the current page. enumerate Extended version of the enumerate environment. fontsmpl Package and test file for producing ‘font samples’. ftnright Place all footnotes in the right-hand column in two-column mode. indentfirst Indent The first paragraph of sections, etc. layout Show the page layout defined by the current document class. multicol Typeset text in columns, with the length of the columns ‘balanced’. rawfonts Preload fonts using the old internal font names of L^{A}TEX 2.09. See

Section 5.2. somedefs Selective handling of package options. (Used by the rawfonts pack-

age.) showkeys Prints the ‘keys’ used by \label, \ref, \cite etc.; useful whilst

drafting. theorem Flexible declaration of ‘theorem-like’ environments.

varioref ‘Smart’ handling of page references.

verbatim Flexible extension of the verbatim environment.

xr Cross reference other ‘external’ documents.

xspace ‘Smart space’ command that helps you to avoid the common mistake of missing spaces after command names.

3 Commands

This section describes the new commands available in L^{A}TEX 2ε. They are covered in more detail in L^{A}TEX: A Document Preparation System and in The L^{A}TEX Companion.

3.1 Initial commands

Initial commands can appear only before the \documentclass line.

\begin{filecontents} {⟨file-name⟩}

⟨file-contents⟩

\end{filecontents}

The filecontents environment is intended for bundling within a single document file the contents of packages, options, or other files. When the document file is run through L^{A}TEX 2ε the body of this environment is written verbatim (preceded by a comment line) to a file whose name is given as the environment’s only argument. However, if that file already exists then nothing happens except for an information message.

Only normal ASCII text characters (i.e. 7-bit visible text) should be included in a filecontents environment. Anything else, such as tab characters, form-feeds or 8-bit characters, should not be included in a filecontents environment.

Tabs and form feeds produce a warning, explaining that they are turned into spaces or blank lines, respectively. What happens to 8-bit characters depends on the TEX installation and is in general unpredictable.

The filecontents environment is used for including L^{A}TEX files. For other plain text files (such as Encapsulated PostScript files), you should use the filecontents* environment which does not add a comment line.

These environments are allowed only before \documentclass. This ensures that any packages that have been bundled in the document are present when needed.

3.2 Preamble commands

The changes to the preamble commands are intentionally designed to make L^{A}TEX 2ε documents look clearly different from old documents. The commands should be used only before \begin{document}.

\documentclass [⟨option-list⟩] {⟨class-name⟩} [⟨release-date⟩]

This command replaces the L^{A}TEX2.09 command \documentstyle.

There must be exactly one \documentclass command in a document; and it must come after the filecontents environments, if any, but before any other commands.

The ⟨option-list⟩ is a list of options, each of which may modify the formatting of elements which are defined in the ⟨class-name⟩ file, as well as those in all following \usepackage commands (see below).

The optional argument ⟨release-date⟩ can be used to specify the earliest desired release date of the class file; it should contain a date in the format yyyy/mm/dd. If a version of the class older than this date is found, a warning is issued.

For example, to specify a two-column article, using a version of article.cls released after June 1994, you specify:

   \documentclass[twocolumn]{article}[1994/06/01]

\documentstyle [⟨option-list⟩] {⟨class-name⟩}

This command is still supported for compatibility with old files. It is essentially the same as \documentclass except that it invokes L^{A}TEX 2.09 compatibility mode. It also causes any options in the ⟨option-list⟩ that are not processed by the class file to be loaded as packages after the class has been loaded. See Section 4 for more details on L^{A}TEX 2.09 compatibility mode.

\usepackage [⟨option-list⟩] {⟨package-name⟩} [⟨release-date⟩]

Any number of \usepackage commands is allowed. Each package file (as denoted by ⟨package-name⟩) defines new elements (or modifies those defined in the class file loaded by the ⟨class-name⟩ argument of the \documentclass command). A package file thus extends the range of documents which can be processed.

The ⟨option-list⟩ argument can contain a list of options, each of which can modify the formatting of elements which are defined in this ⟨package-name⟩ file.

As above, ⟨release-date⟩ can contain the earliest desired release date of the package file in the format yyyy/mm/dd; if an older version of the package is found, a warning is issued.

For example, to load the graphics package for the dvips driver, using a version of graphics.sty released after June 1994, you write:

   \usepackage[dvips]{graphics}[1994/06/01]

Each package is loaded only once. If the same package is requested more than once, nothing happens in the second or following attempt unless the package has been requested with options that were not given in the original \usepackage. If such extra options are specified then an error message is produced. See Section 5 how to resolve this problem.

As well as processing the options given in the ⟨option-list⟩ of the \usepackage command, each package processes the ⟨option-list⟩ of the \documentclass command as well. This means that any option which should be processed by every package (to be precise, by every package that specifies an action for it) can be specified just once, in the \documentclass command, rather than being repeated for each package that needs it.

\listfiles

If this command is placed in the preamble then a list of the files read in (as a result of processing the document) will be displayed on the terminal (and in the log file) at the end of the run. Where possible, a short description will also be produced.

Warning: this command will list only files which were read using L^{A}TEX commands such as \input{⟨file⟩} or \include{⟨file⟩}. If the file was read using the primitive TEX file input command then it will not be listed.

\setcounter{errorcontextlines} {⟨num⟩}

TEX3 introduced a new primitive \errorcontextlines which controls the format of error messages. L^{A}TEX 2ε provides an interface to this through the standard \setcounter command. As most L^{A}TEX users do not want to see the internal definitions of L^{A}TEX commands each time they make an error, L^{A}TEX 2ε sets this to −1 by default.

3.3 Document structure

The book document class introduces new commands to indicate document structure.

\frontmatter
\mainmatter
\backmatter

These commands indicate the beginning of the front matter (title page, table of contents and prefaces), main matter (main text) and back matter (bibliography, indexes and colophon).

3.4 Definitions

In L^{A}TEX, commands can have both mandatory and optional arguments, for example in:

   \documentclass[11pt]{article}

the 11pt argument is optional, whereas the article class name is mandatory.

In L^{A}TEX 2.09 users could define commands with arguments, but these had to be mandatory arguments. With L^{A}TEX 2ε, users can now define commands and environments which also have one optional argument.

\newcommand {⟨cmd⟩} [⟨num⟩] [⟨default⟩] {⟨definition⟩} \renewcommand {⟨cmd⟩} [⟨num⟩] [⟨default⟩] {⟨definition⟩}

These commands have a new, second, optional argument; this is used for defining commands which themselves take one optional argument. This new argument is best introduced by means of a simple (and hence not very practical) example:

   \newcommand{\example}[2][YYY]{Mandatory arg: #2;
                                 Optional arg: #1.}

This defines \example to be a command with two arguments, referred to as #1 and #2 in the {⟨definition⟩}—nothing new so far. But by adding a second optional argument to this \newcommand (the [YYY]) the first argument (#1) of the newly defined command \example is made optional with its default value being YYY.

Thus the usage of \example is either:

   \example{BBB}

which prints:

Mandatory arg: BBB; Optional arg: YYY. or:

   \example[XXX]{AAA}

which prints:

Mandatory arg: AAA; Optional arg: XXX.

The default value of the optional argument is YYY. This value is specified as the [⟨default⟩] argument of the \newcommand that created \example.

As another more useful example, the definition:

   \newcommand{\seq}[2][n]{\lbrace #2_{0},\ldots,\,#2_{#1} \rbrace}

means that the input $\seq{a}$ produces the formula {a_{0}, …, a_{n}}, whereas the input $\seq[k]{x}$ produces the formula {x_{0}, …, x_{k}}.

In summary, the command:

\newcommand {⟨cmd⟩} [⟨num⟩] [⟨default⟩] {⟨definition⟩}

defines ⟨cmd⟩ to be a command with ⟨num⟩ arguments, the first of which is optional and has default value ⟨default⟩.

Note that there can only be one optional argument but, as before, there can be up to nine arguments in total.

\newenvironment {⟨cmd⟩} [⟨num⟩] [⟨default⟩] {⟨beg-def ⟩} {⟨end-def ⟩} \renewenvironment {⟨cmd⟩} [⟨num⟩] [⟨default⟩] {⟨beg-def ⟩} {⟨end-def ⟩}

L^{A}TEX 2ε also supports the creation of environments that have one optional argument. Thus the syntax of these two commands has been extended in the same way as that of \newcommand.

\providecommand {⟨cmd⟩} [⟨num⟩] [⟨default⟩] {⟨definition⟩}

This takes the same arguments as \newcommand. If ⟨cmd⟩ is already defined then the existing definition is kept; but if it is currently undefined then the effect of \providecommand is to define ⟨cmd⟩ just as if \newcommand had been used.

All the above five ‘defining commands’ now have *-forms that are usually the

New feature 1994/12/01

better form to use when defining commands with arguments, unless any of these arguments is intended to contain whole paragraphs of text. Moreover, if you ever do find yourself needing to use the non-star form then you should ask whether that argument would not better be treated as the contents of a suitably defined environment.

3.5 Boxes

These next three commands for making LR-boxes all existed in L^{A}TEX 2.09. They have been enhanced in two ways.

\makebox [⟨width⟩] [⟨pos⟩] {⟨text⟩} \framebox [⟨width⟩] [⟨pos⟩] {⟨text⟩} \savebox {⟨cmd⟩} [⟨width⟩] [⟨pos⟩] {⟨text⟩}

One small but far-reaching change for L^{A}TEX 2ε is that, within the ⟨width⟩ argument only, four special lengths can be used. These are all dimensions of the box that would be produced by using simply \mbox{⟨text⟩}:

\height its height above the baseline;

\depth its depth below the baseline;

     \totalheight the sum of \height and \depth;

\width its width.

Thus, to put ‘hello’ in the centre of a box of twice its natural width, you would use:

   \makebox[2\width]{hello}

Or you could put f into a square box, like this: f

   \framebox{\makebox[\totalheight]{\itshape f\/}}

Note that it is the total width of the framed box, including the frame, which is

set to \totalheight.

The other change is a new possibility for ⟨pos⟩: s has been added to l and r. If ⟨pos⟩ is s then the text is stretched the full length of the box, making use of any ‘rubber lengths’ (including any inter-word spaces) in the box’s contents. If no such ‘rubber length’ is present, an ‘underfull box’ will probably be produced.

\parbox [⟨pos⟩] [⟨height⟩] [⟨inner-pos⟩] {⟨width⟩} {⟨text⟩} \begin{minipage} [⟨pos⟩] [⟨height⟩] [⟨inner-pos⟩] {⟨width⟩}

⟨text⟩

\end{minipage}

As for the box commands above, \height, \width, etc. may be used in the [⟨height⟩] argument to denote the natural dimensions of the box.

The ⟨inner-pos⟩ argument is new in L^{A}TEX 2ε. It is the vertical equivalent to the ⟨pos⟩ argument for \makebox, etc, determining the position of ⟨text⟩ within the box. The ⟨inner-pos⟩ may be any one of t, b, c, or s, denoting top, bottom, centred, or ‘stretched’ alignment respectively. When the ⟨inner-pos⟩ argument is not specified, L^{A}TEX gives it same value as ⟨pos⟩ (this could be the latter’s default value).

\begin{lrbox} {⟨cmd⟩}

⟨text⟩

\end{lrbox}

This is an environment which does not directly print anything. Its effect is to save the typeset ⟨text⟩ in the bin ⟨cmd⟩. Thus it is like \sbox{⟨cmd⟩}{⟨text⟩}, except that any white space before or after the contents ⟨text⟩ is ignored.

This is very useful as it enables both the \verb command and the verbatim environment to be used within ⟨text⟩.

It also makes it possible to define, for example, a ‘framed box’ environment. This is done by first using this environment to save some text in a bin ⟨cmd⟩ and then calling \fbox{\usebox{⟨cmd⟩}}.

The following example defines an environment, called fmpage, that is a framed version of minipage.

   \newsavebox{\fmbox}
   \newenvironment{fmpage}[1]
     {\begin{lrbox}{\fmbox}\begin{minipage}{#1}}
     {\end{minipage}\end{lrbox}\fbox{\usebox{\fmbox}}}

3.6 Measuring things

The first of these next commands was in L^{A}TEX 2.09. The two new commands are the obvious analogues.

\settowidth {⟨length-cmd⟩} {⟨lr text⟩} \settoheight {⟨length-cmd⟩} {⟨lr text⟩} \settodepth {⟨length-cmd⟩} {⟨lr text⟩}

3.7 Line endings

The command \\, which is used to indicate a line-end in various places, is now New description a robust command when used within arguments such as section titles. 1994/12/01 Also, because it is often necessary to distinguish which type of line is to be ended, we have introduced the following new command; it has the same argument syntax as that of \\.

\tabularnewline [⟨vertical-space⟩]

New feature 1994/12/01

One example of its use is when the text in the last column of a tabular environment is set with \raggedright; then \tabularnewline can be used to indicate the end of a row of the tabular, whilst \\ will indicate the end of a line of text in a paragraph within the column. This command can be used in the array environment as well as tabular, and also the extended versions of these environments offered by the array and longtable packages in the tools collection.

3.8 Controlling page breaks

Sometimes it is necessary, for a final version of a document, to ‘help’ L^{A}TEX break the pages in the best way. L^{A}TEX 2.09 had a variety of commands for this situation: \clearpage, \pagebreak etc. L^{A}TEX 2ε provides, in addition, commands which can produce longer pages as well as shorter ones.

\enlargethispage {⟨size⟩}
\enlargethispage* {⟨size⟩}

These commands increase the height of a page (from its normal value of \textheight) by the specified amount ⟨size⟩, a rigid length. This change affects only the current page.

This can be used, for example, to allow an extra line to be fitted onto the page or, with a negative length, to produce a page shorter than normal.

The star form also shrinks any vertical white space on the page as much as possible, so as to fit the maximum amount of text on the page.

3.9 Floats

There is a new command, \suppressfloats, and a new ‘float specifier’. These will enable people to gain better control of L^{A}TEX’s float placement algorithm.

\suppressfloats [⟨placement⟩]

This command stops any further floating environments from being placed on the current page. With an optional argument, which should be either t or b (not both), this restriction applies only to putting further floats at the top or at the bottom. Any floats which would normally be placed on this page are placed on the next page instead.

The extra float location specifier: !

This can be used, along with at least one of h, t, b and p, in the location optional argument of a float.

If a ! is present then, just for this particular float, whenever it is processed by the float mechanism the following are ignored:

• all restrictions on the number of floats which can appear;

• all explicit restrictions on the amount of space on a text page which may be occupied by floats or must be occupied by text.

The mechanism will, however, still attempt to ensure that pages are not overfull and that floats of the same type are printed in the correct order.

Note that its presence has no effect on the production of float pages.

A ! specifier overrides the effect of any \suppressfloats command for this particular float.

3.10 Font changing: text

L^{A}TEX 2ε has a very different font selection scheme to L^{A}TEX 2.09. In this section, we give a brief description of the new commands. A more detailed description with examples is given in The L^{A}TEX Companion, and the interface for classand package-writers is described in L^{A}TEX 2ε Font Selection.

\rmfamily
\sffamily
\ttfamily
\mdseries
\bfseries
\upshape
\itshape
\slshape
\scshape

These are font commands whose use is the same as the commands \rm, \bf, etc. The difference is that each command changes just one attribute of the font (the attribute changed is part of the name). One result of this is that, for example, \bfseries\itshape produces both a change of series and a change of shape, to give a bold italic font.

\textrm{⟨text⟩}
\textsf{⟨text⟩}
\texttt{⟨text⟩}
\textmd{⟨text⟩}
\textbf{⟨text⟩}
\textup{⟨text⟩}
\textit{⟨text⟩}
\textsl{⟨text⟩}
\textsc{⟨text⟩}
\emph{⟨text⟩}

These are one-argument commands; they take as an argument the text which is to be typeset in the particular font. They also automatically insert italic corrections where appropriate; if you do not like the result, you can add an italic correction with \/ or remove it with \nocorr. The \nocorr should always be the first or last thing within the {⟨text⟩} argument.

3.11 Font changing: math

Most of the fonts used within math mode do not need to be explicitly invoked; but to use letters from a range of fonts, the following class of commands is provided.

\mathrm {⟨letters⟩}
\mathnormal {⟨letters⟩}
\mathcal {⟨letters⟩}
\mathbf {⟨letters⟩}
\mathsf {⟨letters⟩}
\mathtt {⟨letters⟩}
\mathit {⟨letters⟩}

These are also one-argument commands which take as an argument the letters which are to be typeset in the particular font. The argument is processed in math mode so spaces within it will be ignored. Only letters, digits and accents have their font changed, for example $\mathbf{\tilde A \times 1}$ produces Ã × 1.

3.12 Ensuring math mode

\ensuremath {⟨math commands⟩}

In L^{A}TEX 2.09, if you wanted a command to work both in math mode and in text mode, the suggested method was to define something like:

   \newcommand{\Gp}{\mbox{$G_p$}}

Unfortunately, the \mbox stops \Gp changing size correctly in (for instance) subscripts or a fraction.

In L^{A}TEX 2ε you can define it thus:

   \newcommand{\Gp}{\ensuremath{G_p}}

Now \Gp will work correctly in all contexts.

This is because the \ensuremath does nothing, producing simply G_p, when \Gp is used within math mode; but it ensures that math mode is entered (and exited) as required when \Gp is used in text mode.

3.13 Text commands: all encodings

One of the main differences between L^{A}TEX 2ε and L^{A}TEX 2.09 is that L^{A}TEX 2ε New description can deal with fonts in arbitrary encodings. (A font encoding is the sequence 1994/12/01 of characters in the font—for example a Cyrillic font would have a different encoding than a Greek font.)

The main two encodings for Latin languages such as English or German are OT1 (Donald Knuth’s 7-bit encoding, which has been used during most of TEX’s lifetime) and T1 (the new 8-bit ‘Cork’ encoding).

L^{A}TEX 2.09 only supported the OT1 encoding, whereas L^{A}TEX 2ε has support for both OT1 and T1 built-in. The next section will cover the new commands which are available if you have T1-encoded fonts. This section describes new commands which are available in all encodings.

Most of these commands provide characters which were available in L^{A}TEX 2.09 already. For example \textemdash gives an ‘em dash’, which was available in L^{A}TEX 2.09 by typing ---. However, some fonts (for example a Greek font) may not have the --- ligature, but you will still be able to access an em dash by

typing \textemdash.

\r{⟨text⟩} New feature 1994/12/01

This command gives a ‘ring’ accent, for example ‘o̊’ can be typed \r{o}.

\SS

New feature 1994/12/01

This command produces a German ‘SS’, that is a capital ‘ß’. This letter can hyphenate differently from ‘SS’, so is needed for entering all-caps German.

New feature 1994/12/01

\textcircled{⟨text⟩}

This command is used to build ‘circled characters’ such as \copyright. For example \textcircled\{a\} produces \begin{matrix}a \\ ◯ \end{matrix}.

\textcompwordmark

New feature 1994/12/01

This command is used to separate letters which would normally ligature. For example ‘fi’ is produced with f\textcompwordmark i. Note that the ‘f’ and ‘i’ have not ligatured to produce ‘fi’. This is rarely useful in English (‘shelfful’ is a rare example of where it might be used) but is used in languages such as German.

\textvisiblespace

New feature 1994/12/01

This command produces a ‘visible space’ character ‘ ’. This is sometimes used in computer listings, for example ‘type hello world’.

\textemdash \textendash \textexclamdown \textquestiondown
\textquotedblleft \textquotedblright \textquoteleft \textquoteright

New feature 1994/12/01

These commands produce characters which would otherwise be accessed via ligatures:

ligature character command

                ---  —  \textemdash
                 --  –  \textendash
                 !`  ¡  \textexclamdown
                 ?`  ¿  \textquestiondown
                 ``  “  \textquotedblleft
                 ''  ”  \textquotedblright
                 `  ‘  \textquoteleft
                 '  ’  \textquoteright

The reason for making these characters directly accessible is so that they will work in encodings which do not have these characters.

\textbullet \textperiodcentered

New feature 1994/12/01

These commands allow access to characters which were previously only available in math mode:

math command character text command

           \bullet  •  \textbullet
           \cdot  ·  \textperiodcentered

3.14 Text commands: the T1 encoding

The OT1 font encoding is fine for typesetting in English, but has problems when New description typesetting other languages. The T1 encoding solves some of these problems, 1994/12/01 by providing extra characters (such as ‘eth’ and ‘thorn’), and it allows words containing accented letters to be hyphenated (as long as you have a package like babel which allows for non-American hyphenation).

This section describes the commands you can use if you have the T1 fonts. To use them, you need to get the DC fonts, or the T1-encoded PostScript fonts, as used by psnfss. All these fonts are available by anonymous ftp in the Comprehensive TEX Archive, and are also available on CD-ROM in 4all TEX and Prime Time Freeware (both available from the TEX Users Group).

You can then select the T1 fonts by saying:

   \usepackage[T1]{fontenc}

This will allow you to use the commands in this section.

Note: Since this document must be processable on any site running an upto-date L^{A}TEX, it does not contain any characters that are present only in T1encoded fonts. This means that this document cannot show you what these glyphs look like! If you want to see them, then run L^{A}TEX on the document fontsmpl and respond ‘cmr’ when it prompts you for a family name.

\k{⟨text⟩} New feature 1994/12/01

This command produces an ‘ogonek’ accent.

\DH \DJ \NG \TH \dh \dj \ng \th

New feature 1994/12/01

These commands produce characters ‘eth’, ‘dbar’, ‘eng’, and ‘thorn’.

\guillemotleft \guillemotright \guilsinglleft \guilsinglright
\quotedblbase \quotesinglbase \textquotedbl

New feature 1994/12/01

These commands produce various sorts of quotation mark. Rough representations of them are: ≪a≫ <a> ,,a” ,a’ and "a".

3.15 Logos

\LaTeX \LaTeXe

\LaTeX (producing ‘L^{A}TEX’) is still the ‘main’ logo command, but if you need to refer to the new features, you can write \LaTeXe (producing ‘L^{A}TEX 2ε’).

3.16 Picture commands

\qbezier[⟨N ⟩](⟨AX ⟩,⟨AY ⟩)(⟨BX ⟩,⟨BY ⟩)(⟨CX ⟩,⟨CY ⟩) \bezier{⟨N ⟩}(⟨AX ⟩,⟨AY ⟩)(⟨BX ⟩,⟨BY ⟩)(⟨CX ⟩,⟨CY ⟩)

The \qbezier command can be used in picture mode to draw a quadratic Bezier curve from position (⟨AX ⟩,⟨AY ⟩) to (⟨CX ⟩,⟨CY ⟩) with control point (⟨BX ⟩,⟨BY ⟩). The optional argument ⟨N ⟩ gives the number of points on the curve.

For example, the diagram:

[figure]

is drawn with:

   \begin{picture}(50,50)
      \thicklines
      \qbezier(0,0)(0,50)(50,50)
      \qbezier[20](0,0)(50,0)(50,50)
      \thinlines
      \put(0,0){\line(1,1){50}}
   \end{picture}

The \bezier command is the same, except that the argument ⟨N ⟩ is not optional. It is provided for compatibility with the L^{A}TEX 2.09 bezier document style option.

3.17 Old commands

\samepage

The \samepage command still exists but is no longer being maintained. This is because it only ever worked erratically; it does not guarantee that there will be no page-breaks within its scope; and it can cause footnotes and marginals to be wrongly placed.

We recommend using \enlargethispage in conjunction with page-break commands such as \newpage and \pagebreak to help control page breaks.

\SLiTeX

Since SliTEX no longer exists, the logo is no longer defined in the L^{A}TEX kernel. A suitable replacement is \textsc{Sli\TeX}. The SliTEX logo is defined in L^{A}TEX 2.09 compatibility mode.

\mho \Join \Box \Diamond \leadsto
\sqsubset \sqsupset \lhd \unlhd \rhd \unrhd

These symbols are contained in the L^{A}TEX symbol font, which was automatically loaded by L^{A}TEX 2.09. However, TEX has room for only sixteen math font families; thus many users discovered that they ran out. Because of this, L^{A}TEX does not load the L^{A}TEX symbol font unless you use the latexsym package.

These symbols are also made available, using different fonts, by the amsfonts package, which also defines a large number of other symbols. It is supplied by the American Mathematical Society.

The latexsym package is loaded automatically in L^{A}TEX 2.09 compatibility mode.

4 L^{A}TEX 2.09 documents

L^{A}TEX 2ε can process (almost) any L^{A}TEX 2.09 document, by entering L^{A}TEX 2.09 compatibility mode. Nothing has changed, you run L^{A}TEX in the same way you always did, and you will get much the same results.

The reason for the ‘almost’ is that some L^{A}TEX 2.09 packages made use of lowlevel unsupported features of L^{A}TEX. If you discover such a package, you should find out if it has been updated to work with L^{A}TEX 2ε. Most packages will still work with L^{A}TEX 2ε—the easiest way to find out whether a package still works is to try it!

L^{A}TEX 2.09 compatibility mode is a comprehensive emulation of L^{A}TEX 2.09, but at the cost of time. Documents can run up to 50% slower in compatibility mode than they did under L^{A}TEX 2.09. In addition, many of the new features of L^{A}TEX 2ε are not available in L^{A}TEX 2.09 compatibility mode.

4.1 Font selection problems

When using compatibility mode, it is possible that you will find problems with font-changing commands in some old documents. These problems are of two types:

• producing error messages;

• not producing the font changes you expected.

In case of error messages it is possible that the document (or an old style file used therein) contains references to old internal commands which are no longer defined, see Section 5.2 for more information if this is the case.

L^{A}TEX 2.09 allowed sites to customize their L^{A}TEX installation, which resulted in documents producing different results on different L^{A}TEX installations. L^{A}TEX 2ε no longer allows so much customization but, for compatibility with old documents, the local configuration file latex209.cfg is loaded every time L^{A}TEX 2ε enters L^{A}TEX 2.09 compatibility mode.

For example, if your site was customized to use the New Font Selection Scheme (NFSS) with the oldlfont option, then you can make L^{A}TEX 2ε emulate this by creating a latex209.cfg file containing the commands:

\ExecuteOptions{oldlfont}\RequirePackage{oldlfont}

Similarly, to emulate NFSS with the newlfont option, you can create a latex209.cfg file containing:

\ExecuteOptions{newlfont}\RequirePackage{newlfont}

4.2 Native mode

To run an old document faster, and use the new features of L^{A}TEX 2ε, you should try using L^{A}TEX 2ε native mode. This is done by replacing the command:

\documentstyle[⟨options⟩,⟨packages⟩]{⟨class⟩}

with:

     \documentclass[⟨options⟩]{⟨class⟩}
     \usepackage{latexsym,⟨packages⟩}

However, some documents which can be processed in L^{A}TEX 2.09 compatibility mode may not work in native mode. Some L^{A}TEX 2.09 packages will only work with L^{A}TEX 2ε in 2.09 compatibility mode. Some documents will cause errors because of L^{A}TEX 2ε’s improved error detection abilities.

But most L^{A}TEX 2.09 documents can be processed by L^{A}TEX 2ε’s native mode with the above change. Again, the easiest way to find out whether your documents can be processed in native mode is to try it!

5 Problems

This section describes some of the things which may go wrong when using L^{A}TEX 2ε, and what you can do about it.

5.1 New error messages

L^{A}TEX 2ε has a number of new error messages. Please also note that many error messages now produce further helpful information if you press h in response to the error prompt.

Option clash for package ⟨package⟩.

The named package has been loaded twice with different options. If you enter h you will be told what the options were, for example, if your document contained:

   \usepackage[foo]{fred}
   \usepackage[baz]{fred}

then you will get the error message:

   Option clash for package fred.

and typing h at the ? prompt will give you:

   The package fred has already been loaded with options:
     [foo]
   There has now been an attempt to load it with options:
     [baz]
   Adding the line:
     \usepackage[foo,baz]{fred}
   to your document may fix this.
   Try typing <return> to proceed.

The cure is, as suggested, to load the package with both sets of options. Note that since L^{A}TEX packages can call other packages, it is possible to get a package option clash without explicitly requesting the same package twice.

Command ⟨command⟩ not provided in base NFSS.

The ⟨command⟩ is not provided by default in L^{A}TEX 2ε. This error is generated by using one of the commands:

   \mho \Join \Box \Diamond \leadsto
   \sqsubset \sqsupset \lhd \unlhd \rhd \unrhd

which are now part of the latexsym package. The cure is to add:

   \usepackage{latexsym}

in the preamble of your document.

LaTeX2e command ⟨command⟩ in LaTeX 2.09 document.

The ⟨command⟩ is a L^{A}TEX 2ε command but this is a L^{A}TEX 2.09 document. The cure is to replace the command by a L^{A}TEX 2.09 command, or to run document in native mode, as described in Section 4.2.

NFSS release 1 command \newmathalphabet found.

The command \newmathalphabet was used by the New Font Selection Scheme Release 1 but it has now been replaced by \DeclareMathAlphabet, the use of which is described in L^{A}TEX 2ε Font Selection.

The best cure is to update the package which contained the \newmathalphabet command. Find out if there is a new release of the package, or (if you wrote the package yourself) consult L^{A}TEX 2ε Font Selection for the new syntax of font commands.

If there is no updated version of the package then you can cure this error by using the newlfont or oldlfont package, which tells L^{A}TEX which version of \newmathalphabet should be emulated.

You should use oldlfont if the document selects math fonts with syntax such as this:

     {\cal A}, etc.

Use newlfont if the document’s syntax is like this:

     \cal{A}, etc.
Text for \verb command ended by end of line.

The \verb command has been begun but not ended on that line. This usually means that you have forgotten to put in the end-character of the \verb command.

Illegal use of \verb command.

The \verb command has been used inside the argument of another command. This has never been allowed in L^{A}TEX—often producing incorrect output without any warning—and so L^{A}TEX 2ε produces an error message.

5.2 Old internal commands

A number of L^{A}TEX 2.09 internal commands have been removed, since their functionality is now provided in a different way. See L^{A}TEX 2ε for Class and Package Writers for more details of the new, supported interface for class and package writers.

\tenrm \elvrm \twlrm . . .
\tenbf \elvbf \twlbf . . .
\tensf \elvsf \twlsf . . .

⋮

These commands provided access to the seventy fonts preloaded by L^{A}TEX 2.09. In contrast, L^{A}TEX 2ε normally preloads at most fourteen fonts, which saves a lot of font memory; but a consequence is that any L^{A}TEX file which used the above commands to directly access fonts will no longer work.

Their use will usually produce an error message such as:

   ! Undefined control sequence.
   l.5 \tenrm

The cure for this is to update the document to use the new font-changing commands provided by L^{A}TEX 2ε; these are described in L^{A}TEX 2ε Font Selection.

If this is not possible then, as a last resort, you can use the rawfonts package, which loads the seventy L^{A}TEX 2.09 fonts and provides direct access to them using the old commands. This takes both time and memory. If you do not wish to load all seventy fonts, you can select some of them by using the only option to rawfonts. For example, to load only tenrm and tenbf you write:

   \usepackage[only,tenrm,tenbf]{rawfonts}

The rawfonts package is distributed as part of the L^{A}TEX tools software, see Section 2.4.

5.3 Old files

One of the more common mistakes in running L^{A}TEX is to read in old versions of packages instead of the new versions. If you get an incomprehensible error message from a standard package, make sure you are loading the most recent version of the package. You can find out which version of the package has been loaded by looking in the log file for a line like:

   Package: fred 1994/06/01 v0.01 Fred's package.

You can use the ⟨release-date⟩ options to \documentclass and \usepackage to make sure that you are getting a suitably recent copy of the document class or package. This is useful when sending a document to another site, which may have out-of-date software.

5.4 Where to go for more help

If you can’t find the answer for your problem here, try looking in L^{A}TEX: A Document Preparation System or The L^{A}TEX Companion. If you have a problem with installing L^{A}TEX, look in the installation guide files which come with the distribution.

If this doesn’t help, contact your local L^{A}TEX guru or local L^{A}TEX mailing list.

If you think you’ve discovered a bug then please report it! First, you should find out if the problem is with a third-party package or class. If the problem is caused by a package or class other than those listed in Section 2 then please report the problem to the author of the package or class, not to the L^{A}TEX3 project team.

If the bug really is with core L^{A}TEX then you should create a short, self-contained document which exhibits the problem. You should run a recent (less than a year old) version of L^{A}TEX on the file and then run L^{A}TEX on latexbug.tex. This will create an error report which you should send, together with the sample document and log file, to the L^{A}TEX bugs address which can be found in the file

latexbug.tex or bugs.txt.

6 Enjoy!

We certainly hope you will enjoy using the new standard L^{A}TEX but, if this is not possible, we hope that you will enjoy success and fulfillment as a result of the documents which it will help you to create.

If you find that the contribution of L^{A}TEX to your life is such that you would like to support the work of the project team, then please read Section 1.2 and discover practical ways to do this.

References

[1] Michel Goossens, Frank Mittelbach and Alexander Samarin. The L^{A}TEX Companion. Addison-Wesley, Reading, Massachusetts, 1994.

[2] Donald E. Knuth. The TEXbook. Addison-Wesley, Reading, Massachusetts, 1986. Revised to cover TEX3, 1991.

[3] Leslie Lamport. L^{A}TEX: A Document Preparation System. Addison-Wesley, Reading, Massachusetts, second edition, 1994.